The area of a rectangle is 36 square inches. The perimeter of the rectangle is 36 inches. What are the dimensions of the rectangle to the nearest hundredth of an inch?
Answer by stanbon(75887)

The area of a rectangle is 36 square inches. The perimeter of the rectangle is 36 inches. What are the dimensions of the rectangle to the nearest hundredth of an inch?
-------
Area:: L*W = 36
-------
Perimeter:: 2(L+W) = 36
L + W = 18
===================
Substitute for "L" and solve for "W"::
(18-W)W = 36
-W^2 + 18W = 36
W^2 - 18W + 36 = 0
--------
W = [18 +-sqrt(18^2-4*1*36)]/(2)
-----
W = [18+-sqrt(180)]/(2)
-----
W = (18+6sqrt(5))]/(2)
W = 15.71 inches
------
Solve for "L"::
L + W = 18
L + 15.71 = 18
L = 2.29 inches
================
Cheers,
Stan H.
=================